Line 2 is perpendicular to a line that passes through the points (4,-3) and (-2,6). How do you solve this using the slope intercept formula (y=mx+b)? I tried to solve it about a hundred times and because it involves a fraction I became entirely confused. Help?! Thanks so much!
Answer by Nate(3500)

Use slope formula:
M = (y - y)/(x - x) = (-3 - 6)/(4 + 2) = (-9)/6 = -3/2
Perpendicular would be the inverse of the opposite, so it the slope of your new line is: M = 2/3
I need more information to make the new equation.